# Axion Icebergs: Clockwork ALPs at hadron colliders

### v1

@@ 3 Clockwork axions at colliders
(b) Now, eschewing the notion of $a_{0}$ as a DM candidate, we assume a larger set of CW scalars, namely $N=40$, while keeping the rest of the parameters the same as in II(a). Then, from eq.(14) it is clear that the light axion mass lowers to a value $m_{a_{0}}\sim 1.5\times 10^{-8}$ eV. The mass splittings in this case fall in the range $\sim 0.5-2.7$ GeV — smaller than what we obtained in benchmark II(a) but not small enough to be completely unresolved at the detector. Consequently, the distribution of signal events, as a function of the diphoton invariant mass is expected to be an undulating one.
@@
Benchmark III: Exploring ALPs heavier than the SM Higgs, we consider $m=150$ GeV, $N=40$ and $f=1600$ GeV. Due to the increase in the SSB scale $f$, the QCD axion mass is now even slightly lower than that in benchmark II(b), namely $m_{a_{0}}\approx 9.7\times 10^{-9}$ eV. As before, Fig.4 shows the diphoton cross-sections for the ALP spectrum as well as the characteristic mass-splittings.
@@
With the model parameters defined, we now simulate the diphoton signal events and compare with the corresponding SM backgrounds for the $13$ TeV LHC. We write the Feynman rules of the models in Feynrules Alloul:2013bka and generate UFO files that are then imported in the MadGraph5 Alwall:2014hca event generator for a Monte Carlo simulation of the signal. Further, Pythia8 Sjostrand:2014zea ; Bierlich:2022pfr is used for parton showering and hadronization. Final state objects at the detector level are reconstructed using the fast simulation tool Delphes deFavereau:2013fsa . Within Delphes we choose to simulate the detector effects using the ATLAS detector card for the entirety our study.
@@
4 Signal and Background profiles
While we would be concentrating on ALPs being produced in gluon fusion and decaying into a diphoton pair, note that the exact final state would be dependent on the average mass of the ALPs. In particular, a very light ALP, produced sans any accompanying high-$p_{T}$ particle, would result in a pair of relatively soft photons that would, typically, fail trigger requirements. In other words, a minimal number of high-$p_{T}$ entities must be present, and thus, both signal and background estimations would need to be done accordingly.
4.1 Light ALPs
For Benchmark Point I, the ALP masses are in the $10-30\,\mbox{GeV}$ range, and the final state photons need to be boosted to overcome the detector’s trigger level energy threshold. To achieve this, we follow the strategy employed in the ATLAS search ATLAS:2022abz (and also ref.Mariotti:2017vtv ) for low mass diphoton resonances and consider a diphoton final state with upto two additional jets. To simulate the signal events we first generate [^9] Understandably, this is overwhelmingly dominated by the subprocesses wherein the final state partons are gluons. $pp\to a_{n}+0/1/2\,{\rm partons}$ at the leading order (LO), using MadGraph5, for $\sqrt{s}=13$ TeV with the NNPDF2.3LO PDF set, followed by the two-body decay of the on-shell ALPs [^10] Given that the ALPs are spinless, there are no non-trivial spin correlations. Furthermore, given that they are very narrow, there is no real loss of information or accuracy in the neglect of possible off-shell effects, as can also be confirmed by the use of tools such as the Madspin Artoisenet:2012st module.. This is followed by parton showering and hadronization via Pythia8 Sjostrand:2014zea ; Bierlich:2022pfr . Matching and merging of the matrix elements (including the avoidance of overcounting) for showering is automated using the MLM matching scheme.
@@ 4.1 Light ALPs
Jets: Within Delphes, jets are reconstructed using the FastJet package Cacciari:2011ma . Jet clustering is performed using the anti-$k_{T}$ algorithm Cacciari:2008gp with the jet cone radius parameter $R$ chosen to be 0.4. To be consonant with the ATLAS analyses, we require the minimum transverse momentum of jets $p_{Tj}$ to be 20 GeV and its pseudorapidity to satisfy $|\eta_{j}|<2.5$. Finally, any two jets must be separated by $\Delta R(j,j)>0.7$ where $(\Delta R)\equiv\left[(\Delta\eta)^{2}+(\Delta\phi)^{2}\right]^{1/2}$ is the separation in the pseudorapidity-azimuthal angle plane.
Photons: For identification and isolation of photons, a cone size of $\Delta R=0.2$ around the photon candidate is considered. Denoting the ratio of the sum of transverse momenta of isolated objects (tracks, calorimeter towers, etc) to the candidate’s transverse momentum as $p_{T}^{\text{ratio}}$, it is demanded that $p_{T}^{\text{ratio}}<0.05$. Also incorporated is a $p_{T}$-dependent photon identification efficiency (following the ATLAS analysis of ref.ATLAS:2022abz ) that ranges from $\approx 70\%$ at $p_{T}=22$ GeV to $\approx 90\%$ for $p_{T}>50$ GeV.
In addition, we ensure angular separation between photons or jets by demanding that $\Delta R(\gamma,\gamma)>0.2$ and $\Delta R(\gamma,j)>0.4$. Finally we demand events with at least two photons and at least a jet in the final state by imposing $N_{\gamma}\geq 2$ and $N_{j}\geq 1$. The above conditions are summarised in Table 1 for quick reference.
Channel Acceptance Cuts
@@
$pp\rightarrow a_{n}+0/1/2\,\text{jets}$ $\Delta R=0.2$, $p_{T}>0.5$ GeV, $p_{T}^{\text{ratio}}(\gamma)<0.05$
$a_{n}\rightarrow\gamma\gamma$ Jet identification
@@
Table 1: Acceptance cuts for the final state objects in benchmarks I and II ATLAS:2022abz .
@@
As for the QCD corrections to the process, within the narrow width approximation, we may factorise these separately for the production and the subsequent decay. For such masses of the ALPs, the QCD corrections to the production cross section can largely be summarised in terms of $K$-factors, with $K^{1}_{\sigma}\sim K^{2}_{\sigma}\simeq 2$ Gershtein:2020mwi where the superscript indicates the number of jets in the final state. As for the decay, while the correction to the photonic branching fraction is small, that to the gluonic one is substantial, viz. $K_{gg}=2.1$ Djouadi:2005gi , and serves to scale the diphoton branching fraction (since $\mbox{BR}_{\gamma\gamma}\approx\Gamma_{\gamma\gamma}/\Gamma_{gg}$). For the effective $K$-factor, then, $K^{1,2}_{\sigma}/K_{gg}\approx 1$ for the two benchmark points with light ALPs.
@@
Using MadAnalysis5 Conte:2012fm to analyze the signal events following the acceptance cuts, we order the photons and the jets (wherever applicable) in terms of their $p_{T}$ and present, in Fig.5 (left), the corresponding distributions, for an integrated luminosity of $\mathcal{L}=138\,\mbox{fb}^{-1}$. In this, we use a bin size of $1$ GeV, as in the ATLAS analysis ATLAS:2022abz . While the sharp edges (at 20 GeV) in the jet $p_{T}$  [^11] Note that the events with 2 jets are only a subset of the accepted set of events. distributions are but reflections of the acceptance cut imposed, the fast fall off at higher $p_{T}$s is characteristic of QCD radiation. The positions of the peaks are just caused by an interplay of the two effects. As for the photons, the decay of an ALP, in its rest-frame, is isotropic with each daughter photon having an energy equalling half the ALP mass. This distribution, of course, has to be convoluted with the $p_{T}$ of the ALP itself (as counterbalancing the jet $p_{T}$s). And, finally, there are additional convolutions with both the ALP mass-spectrum and the corresponding production cross sections.
Of particular interest is the diphoton invariant mass distribution (Fig.5 (right)) which encompasses the contributions of the entire spectrum of the $28$ ALPs. It is instructive to note the difference with the total production cross section (as in Fig.2(a)) which shows a maximum for $m_{a}\approx 15$ GeV, owing, as we have discussed, to a combination of the variation in the coupling strength as well as the gluon-gluon flux. However, once the extra jets are required to be radiated, the kinematics (and the mass-dependence of the flux) does change considerably. Even more importantly, the requirements [^12] Further corrections arise from the $p_{T}$-dependence of the detector efficiencies. on $p_{T}(\gamma)$ serve to suppress the contribution from the very low mass ALPs, resulting in the significantly shifted maximum (now around $m_{\gamma\gamma}\approx 20$ GeV) as in Fig.5 (right).
@@
Figure 5: BP-I: event distributions, after applying acceptance cuts alone, vide. Table 1. (left) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (right) the two-photon invariant mass.
@@
For the background, diphoton (associated with additional jets) initiated by $q\bar{q}$– as well as $gg$- within SM are considered. However, since there is a finite probability for a putative jet masquerading as a photon in the detector (this is especially true of a low energy jet depositing a large fraction of its energy in the electromagnetic calorimeter), one needs to consider such contributions as well. Thus, the backgrounds emanate primarily from
@@
Other backgrounds emanate from $e^{\pm}$ not leaving discernible tracks, thereby faking photons. These, though, contribute negligibly and can, therefore, be ignored. For an estimate of the relevant background distribution in the diphoton spectra we again refer to the ATLAS analysis ATLAS:2022abz wherein the fiducial region in the phase-space is defined using the selection cuts on $E^{\gamma}_{T}$, $p^{\gamma\gamma}_{T}$ and $\eta_{\gamma}$ as mentioned in Table 2. The diphoton background [^13] The background distribution as estimated in the ATLAS analysis ATLAS:2022abz for a bin size of 1 GeV is available at https://atlas.web.cern.ch/Atlas/GROUPS/PHYSICS/PAPERS/HIGG-2019-23/. thus determined in the analysis has an uncertainty which is predominantly statistical.
@@
$pp\rightarrow a_{n}+0/1/2\,\text{jets}$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.521.37<|\eta_{\gamma}|<$ 1.52),
@@
Table 2: Selection cuts applied to form the fiducial signal regions for benchmarks I and II ATLAS:2022abz .
To examine the viability of the diphoton signal in the light of the ATLAS analysis, we must subject the former to the same selection criteria as the latter. These are listed in Table 2. As an examination of Fig.5 (left) shows, the $E_{T}(\gamma)$ cut, especially when applied to the subleading photon, does eliminate a non-negligible fraction of the signal events. Even harder is the cut on $p_{T}(\gamma\gamma)$, for this requires that the ALP (and, hence, the radiated off jets) must carry a substantial $p_{T}$. The consequent reduction of events is reflected in the $m_{\gamma\gamma}$ distribution as shown in Fig.6.
@@
A striking aspect to note here (this was already evident in Fig.5 (right)) is that although the diphoton distribution receives contributions from all the ALPs in the spectrum, the separation between the individual resonances is smeared by the detector resolution [^14] We have made use of the energy resolution function for the ATLAS detector’s ECAL AHARROUCHE2006601 as implemented within Delphes. For the phase-space region of interest throughout the presented analysis (i.e. for BP-I, II and III), the energy resolution of the individual photons varies roughly in the range $\sim[0.32,2.3]$ GeV. to a degree that the entire distribution appears as a single broad resonance — an iceberg of axions, so to say — with a FWHM $\sim 16$ GeV. This feature, as previously mentioned, is peculiar to the clockwork spectrum for any mass scale $m$ as long as $N$ is adequately large.
@@
We can now estimate the significance of the signal events over the background bin-by-bin using $S=N_{sig}/\sqrt{N_{sig}+N_{bkg}}$, where $N_{sig}$ and $N_{bkg}$ denote the number of signal and background events, respectively, in a particular bin. Since the background uncertainty is statistics dominated, this gives quite a robust estimate of the sensitivity and can be bettered only by a dedicated search. The bin-wise significances (and the corresponding $p$-values) are depicted in Fig.7, and the maximum individual significance of $S^{1\,\rm GeV}_{138}\sim 1\sigma$ is obtained in the bin corresponding to $m_{\gamma\gamma}\sim 21-22$ GeV. It is tempting to increase the bin widths as this would be expected to substantially increase the per-bin significance owing to the individual peaks being closely packed within the invariant mass range $10-30$ GeV. However, since the existence of a wide band of resonances is characteristic to the clockwork scenario, the shape of the invariant mass distribution could itself behave as a discriminator— more so in the case of small mass-splittings— and by enlarging the bin width one would, obviously, lose information about the underlying profile of the spectrum. In particular, this would help discriminate between a broad resonance and multiple sharp ones [^15] A further discriminant here would be the comparison between the width of the resonance and the total signal size. For a resonance, production cross section is simply related to the total width, whereas the size of the signal is just the product of the cross section, the branching fraction and the overall detector efficiency (the last being a known quantity).. A more useful method of estimating the total signal significance would be to consider the sum $\sum_{i}S_{i}^{2}$ which, of course, is $\chi^{2}$-distributed for the appropriate degrees of freedom. Using this, we find for this benchmark point, the expected cumulative significance to be $S_{c,138}=3.38\sigma$ for $\mathcal{L}=138\,\rm fb^{-1}$. On the same note, it is worth alluding to the prospect that the significance can further build up by a sizable amount with an increase in the luminosity, even by as early as the end of the ongoing Run-3 phase of the LHC with a projected luminosity reach of $\sim 300\,\rm fb^{-1}$. A simple estimate of this can be obtained by scaling the current significance with the luminosity, i.e.
@@
Apropos of the current benchmark point, Fig.8 shows the variation of $S_{c}$ with the SSB scale $f$ for the $138$ $\rm fb^{-1}$ luminosity case alongwith a projection for the $300$ $\rm fb^{-1}$ reach of the LHC. The variation follows the relation $S_{c}\propto f^{-2}$ which can be understood from the fact that for a particular choice of the CW parameters the pertinent cross-sections scale roughly as $f^{-2}$ with the pole masses being quite insensitive to it.
@@
4.2 Intermediate mass ALPs

### v2

@@ III Clockwork axions at colliders
Benchmark II(b): Now, we assume a larger set of CW scalars, namely $N=40$, while keeping the rest of the parameters the same as in II(a). Then, from eq.(15) it is clear that the light axion mass lowers down to a value $m_{a_{0}}\sim 1.5\times 10^{-8}$ eV and the effective decay constant increases to $f_{0}\approx 4\times 10^{14}$ GeV. Note that an enhancement in $f_{0}$ of this magnitude is incompatible with the DM constraint unless $\theta_{0}\ll\mathcal{O}(1)$. Although certain cosmological scenarios could exist where a small initial misalignment angle seems plausible (see ref.Fox et al. (2004) for a discussion), we refrain from delving into those details and, from a strictly phenomenological standpoint of studying different ALP signatures at the colliders, assume a small enough value for $\theta_{0}$ such that the axion relic density is acceptable. The mass splittings in this case fall in the range $\sim 0.5-2.7$ GeV — smaller than what we obtained in benchmark II(a) but not small enough to be completely unresolved at the detector. Consequently, the distribution of signal events, as a function of the diphoton invariant mass is expected to be an undulating one.
Figure 3: Masses and couplings for individual ALPs and corresponding diphoton cross-sections. (top) benchmark II(a) and (bottom) benchmark II(b).
@@
Benchmark III: Exploring ALPs heavier than the SM Higgs, we consider $m=150$ GeV, $N=40$ and $f=1600$ GeV. Due to the increase in the SSB scale $f$, the QCD axion mass is now even slightly lower than that in benchmark II(b), namely $m_{a_{0}}\approx 9.7\times 10^{-9}$ eV with $f_{0}\approx 6\times 10^{14}$ GeV. Naturally, the arguments pertaining to DM abundance as presented for benchmark II(b) are also applicable here. As before, Fig.4 shows the diphoton cross-sections for the ALP spectrum as well as the characteristic mass-splittings.
@@
With the model parameters defined, we now simulate the diphoton signal events and compare with the corresponding SM backgrounds for the $13$ TeV LHC. We write the Feynman rules of the models in Feynrules Alloul et al. (2014) and generate UFO files that are then imported in the MadGraph5 Alwall et al. (2014) event generator for a Monte Carlo simulation of the signal. Further, Pythia8 Sjöstrand et al. (2015); Bierlich et al. (2022) is used for parton showering and hadronization. Final state objects at the detector level are reconstructed using the fast simulation tool Delphes de Favereau et al. (2014). Within Delphes we choose to simulate the detector effects using the ATLAS detector card for the entirety our study.
@@
IV Signal and Background profiles
While we would be concentrating on ALPs being produced in gluon fusion and decaying into a diphoton pair, note that the exact final state would be dependent on their average mass. In particular, a very light ALP, produced sans any accompanying high-$p_{T}$ particle, would result in a pair of relatively soft photons that would, typically, fail trigger requirements. In other words, a minimal number of high-$p_{T}$ entities must be present, and thus, both signal and background estimations would need to be done accordingly.
IV.1 Light ALPs
For Benchmark Point I, the ALP masses are in the $10-30\,\mbox{GeV}$ range, and the final state photons need to be boosted to overcome the detector’s trigger level energy threshold. To achieve this, we follow the strategy employed in the ATLAS search Aad et al. (2023a) (and also ref.Mariotti et al. (2018)) for low mass diphoton resonances and consider a diphoton final state with upto two additional jets. To simulate the signal events we first generate [^18] Understandably, this is overwhelmingly dominated by the subprocesses wherein the final state partons are gluons. $pp\to a_{n}+0/1/2\,{\rm partons}$ at the leading order (LO), using MadGraph5, for $\sqrt{s}=13$ TeV with the NNPDF2.3LO PDF set, followed by the two-body decay of the on-shell ALPs [^19] Given that the ALPs are spinless, there are no non-trivial spin correlations. Furthermore, given that they are very narrow, there is no real loss of information or accuracy in the neglect of possible off-shell effects, as can also be confirmed by the use of tools such as the Madspin Artoisenet et al. (2013) module.. This is followed by parton showering and hadronization via Pythia8 Sjöstrand et al. (2015); Bierlich et al. (2022). Matching and merging of the matrix elements (including the avoidance of overcounting) for showering is automated using the MLM matching scheme.
@@ IV.1 Light ALPs
Jets: Within Delphes, jets are reconstructed using the FastJet package Cacciari et al. (2012). Jet clustering is performed using the anti-$k_{T}$ algorithm Cacciari et al. (2008) with the jet cone radius parameter $R$ chosen to be 0.4. To be consonant with the ATLAS analyses, we require the minimum transverse momentum of jets ($p_{Tj}$) to be 20 GeV and its pseudorapidity to satisfy $|\eta_{j}|<2.5$. Finally, any two jets must be separated by $\Delta R(j,j)>0.7$ where $(\Delta R)\equiv\left[(\Delta\eta)^{2}+(\Delta\phi)^{2}\right]^{1/2}$ is the separation in the pseudorapidity-azimuthal angle plane.
Photons: For identification and isolation of photons, a cone size of $\Delta R=0.2$ around the photon candidate is considered. Denoting the ratio of the sum of transverse momenta of isolated objects (tracks, calorimeter towers, etc) to the candidate’s transverse momentum as $p_{T}^{\text{ratio}}$, it is demanded that $p_{T}^{\text{ratio}}<0.05$. Also incorporated is a $p_{T}$-dependent photon identification efficiency (following the ATLAS analysis of ref.Aad et al. (2023a)) that ranges from $\approx 70\%$ at $p_{T}=22$ GeV to $\approx 90\%$ for $p_{T}>50$ GeV.
In addition, we ensure angular separation between photons or jets by demanding that $\Delta R(\gamma,\gamma)>0.2$ and $\Delta R(\gamma,j)>0.4$. Finally, we demand events with at least two photons and at least a jet in the final state by imposing $N_{\gamma}\geq 2$ and $N_{j}\geq 1$. The above conditions are summarised in Table 1 for quick reference.
Channel : $pp\rightarrow a_{n}+0/1/2\,\text{jets}$   $a_{n}\rightarrow\gamma\gamma$
Acceptance Cuts
@@
$\Delta R=0.2$, $p_{T}>0.5$ GeV, $p_{T}^{\text{ratio}}(\gamma)<0.05$
Jet ident ification
@@
Table 1: Acceptance cuts for the final state objects in benchmarks I and II Aad et al. (2023a).
@@
As for the QCD corrections to the process, within the narrow width approximation, we may factorise these separately for the production and the subsequent decay. For such masses of the ALPs, the QCD corrections to the production cross section can largely be summarised in terms of $K$-factors, with $K^{1}_{\sigma}\sim K^{2}_{\sigma}\simeq 2$ Gershtein et al. (2021) where the superscript indicates the number of jets in the final state. As for the decay, while the correction to the photonic branching fraction is small, that to the gluonic one is substantial, viz. $K_{gg}=2.1$ Djouadi (2008), and serves to scale the diphoton branching fraction (since $\mbox{BR}_{\gamma\gamma}\approx\Gamma_{\gamma\gamma}/\Gamma_{gg}$). For the effective $K$-factor, then, $K^{1,2}_{\sigma}/K_{gg}\approx 1$ for the two benchmark points with light ALPs.
@@
Using MadAnalysis5 Conte et al. (2013) to analyze the signal events following the acceptance cuts, we order the photons and the jets (wherever applicable) in terms of their $p_{T}$ and present, in Fig.5 (top), the corresponding distributions, for an integrated luminosity of $\mathcal{L}=138\,\mbox{fb}^{-1}$. In this, we use a bin size of $1$ GeV, as in the ATLAS analysis Aad et al. (2023a). While the sharp edges (at 20 GeV) in the jet [^20] Note that the events with 2 jets are only a subset of the accepted set of events. $p_{T}$ distributions are but reflections of the acceptance cut imposed, the fast fall off at higher $p_{T}$s is char acteristic of QCD radiation. The positions of the peaks are just caused by an interplay of the two effects. As for the photons, the decay of an ALP, in its rest-frame, is isotropic with each daughter photon having an energy equalling half the ALP mass. This distribution, of course, has to be convoluted with the $p_{T}$ of the ALP itself (as counterbalancing the jet $p_{T}$s). And, finally, there are additional convolutions with both the ALP mass-spectrum and the corresponding production cross sections.
@@
Figure 5: BP-I: event distributions, after applying acceptance cuts alone, vide. Table 1. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
Of particular interest is the diphoton invariant mass distribution (Fig.5 (bottom)) which encompasses the contributions of the entire spectrum of the $28$ ALPs. It is instructive to note the difference with the total production cross section (as in Fig.2 (top)) which shows a maximum for $m_{a}\approx 15$ GeV, owing, as we have discussed, to a combination of the variation in the coupling strength as well as the gluon-gluon flux. However, once the extra jets are required to be radiated, the kinematics (and the mass-dependence of the flux) does change considerably. Even more importantly, the requirements [^21] Further corrections arise from the $p_{T}$-dependence of the detector efficiencies. on $p_{T}(\gamma)$ serve to suppress the contribution from the very low mass ALPs, resulting in the significantly shifted maximum (now around $m_{\gamma\gamma}\approx 20$ GeV) as in Fig.5 (bottom).
@@
For the background, diphoton (associated with additional jets)— initiated by $q\bar{q}$ as well as gggg— within the SM are considered. However, since there is a finite probability for a putative jet masquerading as a photon in the detector (this is especially true of a low energy jet depositing a large fraction of its energy in the electromagnetic calorimeter), one needs to consider such contributions as well. Thus, the backgrounds emanate primarily from
@@
Other backgrounds emanate from $e^{\pm}$ not leaving discernible tracks, thereby faking photons. These, though, contribute negligibly and can, therefore, be largely ignored. For an estimate of the relevant background distribution in the diphoton spectra, we again refer to the ATLAS analysis Aad et al. (2023a) wherein the fiducial region in the phase-space is defined using the selection cuts on $E^{\gamma}_{T}$, $p^{\gamma\gamma}_{T}$ and $\eta_{\gamma}$ as mentioned in Table 2. The diphoton background [^22] The background distribution as estimated in the ATLAS analysis Aad et al. (2023a) for a bin size of 1 GeV is available at https://atlas.web.cern.ch/Atlas/GROUPS/PHYSICS/PAPERS/HIGG-2019-23/. thus determined in the analysis has an uncertainty which is predominantly statistical.
@@
$pp\rightarrow a_{n}+0/1/2\,\text{jets}$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.37<|\eta_{\gamma}|<1.52$),
@@
Table 2: Selection cuts applied to form the fiducial signal regions for benchmarks I and II Aad et al. (2023a).
To examine the viability of the diphoton signal in the light of the ATLAS analysis, we must subject the former to the same selection criteria as the latter. These are listed in Table 2. As an examination of Fig.5 (top) shows, the $E_{T}(\gamma)$ cut, especially when applied to the subleading photon, does eliminate a non-negligible fraction of the signal events. Even harder is the cut on $p_{T}(\gamma\gamma)$, for this requires that the ALP (and, hence, the radiated off jets) must carry a substantial $p_{T}$. The consequent reduction of events is reflected in the $m_{\gamma\gamma}$ distribution as shown in Fig.6.
@@
A striking aspect to note here (this was already evident in Fig.5 (bottom)) is that although the diphoton distribution receives contributions from all the ALPs in the spectrum, the separation between the individual resonances is smeared by the detector resolution [^23] We have made use of the energy resolution function for the ATLAS detector’s ECAL M. Aharrouche, et al, ATLAS Electromagnetic Barrel Calorimeter Collaboration (2006) as implemented within Delphes. For the phase-space region of interest throughout the presented analysis (i.e. for BP-I, II and III), the energy resolution of the individual photons varies roughly in the range $\sim[0.32,2.3]$ GeV. to a degree that the entire distribution appears as a single broad resonance — an iceberg of axions, so to say — with a FWHM $\sim 16$ GeV. This feature, as previously mentioned, is peculiar to the clockwork spectrum for any mass scale $m$ as long as $N$ is adequately large.
@@
We can now estimate the significance of the signal events over the background bin-by-bin using $S=N_{sig}/\sqrt{N_{sig}+N_{bkg}}$, where $N_{sig}$ and $N_{bkg}$ denote the number of signal and background events, respectively, in a particular bin. Since the background uncertainty is statistics dominated, this gives quite a robust estimate of the sensitivity and can be bettered only by a dedicated search. The bin-wise significances (and the corresponding $p$-values) for $\mathcal{L}=138\,\mbox{fb}^{-1}$ with a bin width of $1$ GeV are depicted in Fig.7, and the maximum individual significance of $S^{1\,\rm GeV}_{138}\sim 1\sigma$ is obtained in the bin corresponding to $m_{\gamma\gamma}\sim 21-22$ GeV. It is tempting to increase the bin widths as this would be expected to substantially increase the per-bin significance owing to the individual peaks being closely packed within the invariant mass range 103010-30 GeV. However, since the existence of a wide (and finite) band of resonances is characteristic to the clockwork scenario, the shape of the invariant mass distribution could itself behave as a discriminator— more so in the case of small mass-splittings— and by enlarging the bin width one would, obviously, lose information about the underlying profile of the spectrum. In particular, this would help discriminate between a broad resonance and multiple sharp ones [^24] A further discriminant here would be the comparison between the width of the resonance and the total signal size. For a resonance, production cross section is simply related to the total width, whereas the size of the signal is just the product of the cross section, the branching fraction and the overall detector efficiency (the last being a known quantity).. A more useful method of estimating the total signal significance would be to consider the sum $S_{c}\equiv\sum_{i}S_{i}^{2}$ which, of course, is $\chi^{2}$-distributed for the appropriate degrees of freedom. Using this, we find for this benchmark point, the expected cumulative significance to be $S_{c,138}=3.38\sigma$ for $\mathcal{L}=138\,\rm fb^{-1}$. On the same note, it is worth alluding to the prospect that the significance can further build up by a sizable amount with an increase in the luminosity, even by as early as the end of the ongoing Run-3 phase of the LHC with a projected luminosity reach of $\sim 300\,\rm fb^{-1}$. A simple estimate of this can be obtained by scaling the current significance with the luminosity, i.e.
@@
Apropos of the current benchmark point, Fig.8 shows the variation of $S_{c}$ with the SSB scale $f$ for the $138$ $\rm fb^{-1}$ luminosity case alongwith a projection for the $300$ $\rm fb^{-1}$ reach of the LHC. The variation follows the approximate relation $S_{c}\propto f^{-2}$ which can be understood from the fact that, for a particular choice of the CW parameters, the pertinent cross-sections scale roughly as $f^{-2}$ with the pole masses being quite insensitive to it.
@@
At this stage, it is worth examining if such a signal could have emanated from any other theory. First and foremost, the decay of the resonance into a diphoton pair rules out a spin-one particle. While spin-two or higher are possible, such fundamental particles are disfavoured on theoretical grounds (e.g., a graviton, while well-motivated, would not only be much narrower, but would also decay democratically into other channels). Such arguments narrow down the possibilities to only a scalar or a pseudoscalar. This apart, a width comparable to the mass (as is the case here), would require the couplings to be in the nonperturbative regime Salvio et al. (2016); Franceschini et al. (2016). (For instance, a single ALP resonance with a total decay width and mass $\Gamma(a)\sim m_{a}\sim\mathcal{O}(10{\rm\,GeV})$ would require its decay constant [^25] The inequality takes into account the dependence on the constituents of the anomaly loop. to be $f\lesssim 100$ MeV!). Irrespective of the underlying theory, such a large width would have translated, vide eq.(19), to a total production cross section several orders of magnitude larger than what is the case here. This would, immediately, raise question as to the reason for the suppression of the diphoton branching ratio that such a production cross section (alongwith the observed rates) would imply. Furthermore, an entity that couples so strongly would leave tell-tale signatures in myriad other processes (a trivial example being double diffractive processes leading to a pair of central jets with large rapidity gaps) and, hence, would have been long discovered.
In other words, if such an excess were to be observed, a single resonance hypothesis would be immediately ruled out. If it were to be then interpreted as an overlap of multiple resonances, two issues would need to be considered. The size of the excess imposes restrictions in the plane of the individual total widths and the diphoton branching fraction. On the other hand, a conservative estimate of the number of peaks can be obtained by comparing the width of the excess to the invariant mass resolution operative at such masses. And, finally, the very shape of the excess would dictate that the couplings of the said resonances must conform to a very specific pattern. To reproduce the features in a generic (i.e., non-clockwork) model (or even in multi-axion scenarios involving a new $N$-flavor confining sector Alexander et al. (2023, 2024)) would, thus, require many apparently ad hoc assumptions to be made. On similar grounds, it can also be argued that a broad resonance signal — with a finite bandwidth that is comparable with the characteristic mass scale of the resonance $\sim m$—is incompatible with, say, the Kaluza-Klein spectra of extra-dimensional scenarios consisting, in principle, of an infinite number of states (upto a natural cutoff of the theory), for then the distribution of resonances would extend over a large mass range and truncate only when the cross-sections start to fall near the kinematic limits of the experiment.
It is perhaps worth reiterating at this point that the analysis strategy employed here is complementary to the collider study of the clockwork graviton model carried out in refs.Giudice et al. (2018); Beauchesne and Kats (2020). Of the different strategies investigated therein, a particularly interesting approach involves searching for periodic signals in various kinematic distributions. However, the said technique, while well-suited for the tower-like spectra typically encountered in clockwork models, seems to be more efficient when the resonances are experimentally resolvable. The analysis above, therefore, presents a categorically new perspective to look at clockwork spectra at colliders, especially for the case of light ALPs with tiny mass-splittings.
IV.2 Intermediate mass ALPs